OpenStudy (anonymous):

jane sells 8 times as many Volvos as Melissa. If the difference in the sales is 35, how many cars did Jane sell?

OpenStudy (anonymous):

J is cars sold by Jane, and M car sold by Melissa J = 8M J - M = 35 you can plug 8M from first eq. into second 8M - M = 35 7M = 35 M = 5 J = 8M J = 8(5) J = 40 Jane sold 40 cars, Melissa 5 cars.